Seven inches are equivalent to 17.78 cm. Find the equation converting length in cm to length in inches. What is the length in inches of a 96.2 cm bar?
Found 2 solutions by stanbon, checkley77:
Answer by stanbon(75887)

Seven inches are equivalent to 17.78 cm. Find the equation converting length in cm to length in inches. What is the length in inches of a 96.2 cm bar?
----------------------------
96.2 cm(7 in/17.78 cm) = (96.2*7/17.78) in
---
= 37.87 inches
====================
Cheers,
Stan H.

Answer by checkley77(12844)

17.78/7=2.54 cm. per inch.
96.2/2.54=37.874 in. ans.